OpenStudy (anonymous):

A car is traveling at a rate of 81 kilometers per hour. What is the car's rate in meters per second? How many meters will the car travel in 10 seconds? Do not round your answers.

OpenStudy (anonymous):

?m/s=81km/h * 1h/3600sec * 1000m/1km=81*1000/3600=81/3.6=22.5m/s it will go 225 meters in 10 seconds.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with more?